A bag contain 8 red 6 blue balls.two drawing of each 2 ballasts made find the probability that the first drawing two red balls and second drawing 2 blue balls, if the balls drawn first are replaced before the second draw
Answer by rothauserc(4718)

there are 14 balls in the bag
:
Probability of drawing two red balls = (8/14) * (7/13) = 4/13
:
Probability of drawing two blue balls = (6/14) * (5/13) = 15/91
:
Probability of drawing 2 red balls and 2 blue balls = (4/13) * (15/91) = 60/1183
:
Note I assume a single ball drawn is not replaced until after the second ball is drawn
: